mean relative increase in stock price was about .83%, with a standard deviation of .20%
According to Chebyshev's theorm at least 84% lies between what two numbers.
I understand the theorm is supposed to be 2.5 Standard deviations but I don't see how that number is reached using the Formula. I get cought on the Algebra when the numbers are simplified to the point of
1-(1/K^2) = 0.84
1/k^2 = .16
what is done to make this so,

Answer by stanbon(75887)

According to Chebyshev's theorm at least 84% lies between what two numbers.
I understand the theorm is supposed to be 2.5 Standard deviations but I don't see how that number is reached using the Formula. I get caught on the Algebra when the numbers are simplified to the point of
1-(1/K^2) = 0.84
1/k^2 = .16
k^2 = 1/0.16
k = 1/0.4
k = 2.5
=================
Cheers,
Stan H.